OpenStudy (anonymous):

A jar contains pennies, nickels and dimes. the number of dimes is 1 more than the nickels and the number of pennies is 6 more than the nickels. How many of each can be found in the jar if the total value is $4.80 ?

OpenStudy (lucaz):

$4.80 is the same as 480 cents.. p(1)+n(5)+d(10)=480 dimes=nickels+1 -> d=n+1 pennies=nickels+6 -> p=n+6 (n+6)(1)+n(5)+(n+1)(10)=480 16n=464 -> n=29 d=(29)+1=30 p=(29)+6=35

OpenStudy (anonymous):

Thank lucaz

OpenStudy (lucaz):

no problem